MASS DEPLETION SPECTROSCOPY OF NO_{2} MOLECULE

\underline{DANIIL STOLYAROV}, ELENA POLYAKOVA and CURT WITTIG, Department of Chemistry, University of Southern California, Los Angeles, California 90089.

A mass spectrometric depletion measurement has been performed for the NO_{2} molecule in its bound region as well as above its gas phase dissociation threshold. In the bound region (17700 - 18300 cm^{−1}) that is known to be vibronically chaotic transitions are broadened and blue-shifted relative to their gas-phase counterparts. The spectrum is fitted relatively well by setting all shifts and width equal to 7cm^{−1}. Modest dispersion of these values are consistent with quantum chaos in NO_{2}. Relaxation is dominated by interaction of NO_{2} with its non-superfluid helium nearest neighbors.^{a} In the region above the gas-phase dissociation threshold (25125 - 30000 cm^{−1}) the mass depletion spectrum reveals no fine structure. The fast relaxation due to the interaction with the superfluid helium environment prevents photodecomposition of the NO_{2} molecule over the entire range we explored.

aE. Polyakova, D.Stolyarov, X Zhang, V.V. Kresin and C. Wittig Chem. Phys. Lett. \underline{375} (2003) pp. 253-260